Two cars started from the same point and traveled on a straiht course in opposite directions for exactly 2 hours, at which time they were 208 miles apart. If one car traveled, on average, 8 miles per hour faster than the other car, what was the average speed for each car for the 2-hour trip?

Answer by Alan3354(21605)

Two cars started from the same point and traveled on a straiht course in opposite directions for exactly 2 hours, at which time they were 208 miles apart. If one car traveled, on average, 8 miles per hour faster than the other car, what was the average speed for each car for the 2-hour trip?
-------------
Going in opposite directions, their speeds are added. One car is moving at S mph, the other is S + 8
So S + (S + 8) times 2 = 208
2*(2S + 8) = 208
4S + 16 = 208
4S = 192
S = 48 mph
The faster car moves at 48+8 = 56 mph.